OpenStudy (chrispaul03):

Will Fan And Medal The sum of two numbers is 52. The greater number is 4 more than the smaller number. Which equation can be used to solve for the smaller number? x − (x + 4) = 52 x + (x + 4) = 52 x(x + 4) = 52 x(x − 4) = 52

OpenStudy (anonymous):

can i ask a question

OpenStudy (chrispaul03):

Yes?

OpenStudy (shansha1):

smaller number is (x) greater number is (y) "greater number is 4 more than the smaller number" means y = 4 + x "sum of two numbers is 52" x + y = 52 x + ( 4 + x) = 52 x + (x + 4) = 52